OpenStudy (anonymous):

5(x + 6) <= 3x - 4 My answer is x<=-7 Am I right? :)

OpenStudy (anonymous):

i don't think so, lets check

OpenStudy (anonymous):

\[5x+30\leq 3x-4\] \[2x+30\leq -4\] \[2x\leq -34\]

OpenStudy (anonymous):

i think it is \(x\leq -17\)

OpenStudy (anonymous):

Oh! I forgot the 1 lol, Thank you so much! :)